OpenStudy (anonymous):

integrate (5x^3-3x^2+2x-1)/(x^4+x^2) can i use partial fractions for this?

OpenStudy (slaaibak):

Yes.

OpenStudy (anonymous):

how would i set it up??

OpenStudy (anonymous):

i just have problem setting it up

OpenStudy (mr.math):

\[\frac{5x^3-3x^2+2x-1}{x^2(x^2+1)}=\frac{A}{x}+\frac{B}{x^2}+\frac{Cx+B}{x^2+1}\]

OpenStudy (anonymous):

Partial fractions will haunt you forever so I would master them!